I really need help with this word problem.Thank You.
Radius of a sphere. The radius of a sphere is given by
the formula
r=(0.75v/pi sign)^1/3
where V is its volume. Find the radius of a spherical tank
that has a volume of 32pi sign/3 cubic meters.

Answer by Alan3354(69443)

Radius of a sphere. The radius of a sphere is given by
the formula
r=(0.75v/pi sign)^1/3
where V is its volume. Find the radius of a spherical tank
that has a volume of 32pi sign/3 cubic meters.
-------------
I suppose by "pi sign" you mean pi.
r=(0.75v/pi)^1/3
r = (0.75*(32pi/3)/pi)^(1/3)
Sub for v
r = (0.25*32)^(1/3)
r = 8^(1/3)
r = 2 meters